OpenStudy (anonymous):

input the equation of the line in standard form Ax + By = C the points are (7, -3) and (4, -8)

OpenStudy (anonymous):

First, put these points into a slope-intercept form equation (y=mx+b) Find the slope by y2-y1/x2-x1 which is (-8-(-3))/(4-7)=(-5/-3) or 5/3 Plug that into the slope-intercept equation: y=5/3x+b To find b, plug either one of the points into the equation: -3=5/3(7)+b --> -5 1/3 So the equation is now: y=5/3x-16/3 Put this into standard form: -5/3x+y=-5 1/3 or -5/3x+y=-16/3